OpenStudy (anonymous):

Compute \dfrac{(2/3)^3}{(2/9)^2}.

OpenStudy (phi):

\[ \dfrac{(\frac{2}{3})^3}{(\frac{2}{9})^2} = \frac{\frac{2^3}{3^3}}{\frac{2^2}{9^2}}\] you should multiply top and bottom by the reciprocal of the bottom fraction 9^2/2^2 can you do that ?

OpenStudy (anonymous):

ok

OpenStudy (phi):

I would not multiply anything, just write down the reciprocal times the top and bottom

OpenStudy (phi):

it will look like this \[ \frac{\frac{2^3}{3^3}\cdot \frac{9^2}{2^2}}{\frac{2^2}{9^2} \cdot \frac{9^2}{2^2}} \]

OpenStudy (phi):

notice in the bottom of that mess, the 2^2 in the top divided by 2^2 in the bottom and the 9^2 is divided by 9^2. In other words the bottom is 1 \[ \frac{\frac{2^3}{3^3}\cdot \frac{9^2}{2^2}}{\frac{2^2}{9^2} \cdot \frac{9^2}{2^2}} = \frac{\frac{2^3}{3^3}\cdot \frac{9^2}{2^2}}{1}\] does that part make any sense?

OpenStudy (anonymous):

yep

OpenStudy (phi):

dividing by 1 can be ignored: example 3/1 is 3 or 10/1 is 10, and so on. So we have \[ \frac{2^3}{3^3}\cdot \frac{9^2}{2^2} \]

OpenStudy (phi):

now let's do the twos 2^3 / 2^2 any idea?

OpenStudy (anonymous):

2

OpenStudy (phi):

so now we have \[ \frac{2 \cdot 9^2}{3^3} \]

OpenStudy (anonymous):

ok

OpenStudy (phi):

remember what those little numbers in the upper right mean: multiply by itself so we could do this \[ \frac{ 2 \cdot 9 \cdot 9}{ 3 \cdot 3 \cdot 3} \] can you simplify that ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

what do you get?

OpenStudy (anonymous):

6

OpenStudy (phi):

yes.

OpenStudy (phi):

another way to think of it is \[ \frac{2 \cdot 9^2}{3\cdot 3^2} = \frac{2}{3}\cdot \left(\frac{9}{3}\right)^2 = \frac{2}{3}\cdot 3^2 = 2 \frac{3^2}{3} = 2 \cdot 3 = 6\]

OpenStudy (anonymous):

oh

OpenStudy (phi):

another way is write 9 as 3^2 \[ \frac{2 \cdot 9^2}{3^3} = \frac{2 \cdot \left(3^2\right)^2}{3^3} = \frac{2 \cdot 3^4}{3^3} = 2\cdot \frac{3^4}{3^3} = 2\cdot3=6\]

OpenStudy (anonymous):

ok

OpenStudy (phi):

hopefully some of that makes sense. But it should!

OpenStudy (anonymous):

oh, ok Thanks so much I'm going to be a fan.

OpenStudy (phi):

ty

OpenStudy (phi):

and though you should only do this as a check, the easiest way to solve is type into the google search window (2/3)^3 / (2/9)^2=